A car is travelling at exactly 9 miles per hour and accelerates at a constant rate to exactly 54 miles per hour.

This acceleration takes exactly 16.8 seconds.

What is the rate of acceleration, in miles per hour2?
Answer #1

Concept - convert the time into hour and use equation of kinematics to find the required acceleration as shown below

-------------------------------------------------------------------------------------------------------------------

Let me know in the comment section, if something's wrong
Answer #2

SOLUTION 



Let acceleration be a miles/hr^2 .


Initial speed , v0 = 9 miles/hr


Final speed , vf = 54 miles/hr.


Time taken to reach final speed, t = 16.8 s = 16.8/3600 hr


Now,


vf = v0 + a t 

=> 54 = 9 + a * 16.8/3600

=> a = (54 - 9) / (16.8/3600)

= a = 9642.86 miles/hr^2 (ANSWER).

answered by: Tulsiram Garg